# Exploiting Diffusion Prior for Generalizable Pixel-Level Semantic Prediction

### v1

@@ 4.3 Semantic Segmentation
Semantic segmentation [74] is a fundamental visual understanding task. Since the prediction is categorical, we use a simple conversion for regression models to predict discrete labels. We first generate ground truth labels using the off-the-shelf EVA-2 [15] model. The label maps are then transformed into color maps where each category has different colors. The training and inference of the diffusion model are conducted using the color maps (in the RGB space). During inference, the predicted color maps are converted to categorical label maps by assigning each pixel to its nearest category in the color space.
We report the intersection over union (IoU) and accuracy metrics to measure the in-domain performance in Table 2. The prediction by the proposed scheme is reasonable across all categories compared to the existing methods. Figure 5 demonstrates the results of our method and pseudo ground truth with the in- and out-of-domain (i.e., bedroom images in diverse styles) input images. Particularly in out-of-domain examples, our model gives favorable predictions compared to the off-the-shelf approach, e.g., the painting of the first, the curtain of the second, the window of the third, and the carpet of the fourth images in Figure 5 (b). This validates our idea that leverages the pre-trained T2I diffusion model as the prior for better generalization.
@@ 4.4 Intrinsic Image Decomposition
Intrinsic image decomposition [4] recovers the albedo and shading properties from the RGB image. It facilitates applications such as recoloring [3] and relighting [54]. Similar to the PIE-Net [12] work, we use the mean square error $\mathrm{MSE}$ as the evaluation metric to compute the distance between the predicted and pseudo ground-truth estimations. We showcase the qualitative results in Figure 4 and quantitative comparisons in Table 3. The errors reported for most methods are an order of magnitude larger than ours. Notably, our method and the SPADE approach work well in both in- and out-of-domain settings.
@@
Figure 8: Single vs. multiple diffusion steps. In the single-step approach, we train the U-Net network to directly predict the output from the input image. The single-step approach does not generalize to unseen images well.
@@ 4.5 Ablation Study
Parameterization. We study the effect of employing various parameterizations in this experiment. Specifically, we fine-tune the U-Net model to make different predictions in each reverse diffusion step for obtaining the final prediction: 1) predicting the input image $x$, predicting the output $y$ (similar to $x_{0}$-prediction in standard diffusion models), and v-prediction described in Eq. (3). We formulate each parameterization in detail in the supplementary document. The quantitative results are shown in Table 4 and the qualitative comparisons are presented in Figure 6. Predicting the input image $x$ generates accurate results in the in-domain setting. However, it fails to generalize to unseen domains. On the other hand, predicting the output $y$ demonstrates preferred generalization capability, but tends to produce blurry results with few details. We choose to use the v-prediction approach as it produces accurate results of arbitrary images.
Number of diffusion steps. We analyze the performance of our method with different numbers of diffusion steps in the inference stage. The quantitative comparisons on the normal prediction task are presented in Figure 7. Using $5$ steps strikes a good balance between inference speed and accuracy. Furthermore, we study an extreme case of using a single step, where we train the U-net to directly predict the output $y$ from the input image $x$. Although it shows competitive performance in the in-domain setting according to Figure 7, we find that the performance degrades significantly with the arbitrary input images, as examples shown in Figure 8. Considering the generalizability, inference speed, and accuracy, we use $5$ diffusion steps for all the other experiments.
@@ 5 Conclusion
In this work, we explore to leverage a pre-trained T2I diffusion model for generalizable pixel-level semantic prediction. The core of our method is the design of the deterministic diffusion process that adapts the stochastic T2I framework for the deterministic prediction tasks. With low-rank approximation, the proposed approach learns the target tasks while retaining the inherent generalization ability of the pre-trained T2I model. We show that with only a small number of labeled training data in a limited domain (i.e., $10$K bedroom images), our DMP scheme makes faithful predictions on arbitrary images. We believe that this work establishes a foundation for achieving ultimately-generalizable visual understanding in the future.

### v2

@@ 4.3 Semantic Segmentation
Semantic segmentation [89] is a fundamental visual understanding task. Since the prediction is categorical, we use a simple conversion for regression models to predict discrete labels. We first generate ground-truth labels using the off-the-shelf EVA-2 [17] model. The label maps are then transformed into color maps where each category has different colors. The training and inference of the diffusion model are conducted using the color maps (in the RGB space). During inference, the predicted color maps are converted to categorical label maps by assigning each pixel to its nearest category in the color space.
We report the intersection over union (IoU) and accuracy to measure the in-domain performance in Table 2. The prediction by the proposed scheme is reasonable across all categories compared to the existing methods. Figure 5 demonstrates the results of our method and pseudo ground truth with the in- and out-of-domain (i.e., bedroom images in diverse styles) input images. Particularly in out-of-domain examples, our model gives favorable predictions compared to the off-the-shelf approach, e.g., the painting of the first, the curtain of the second, the window of the third, and the carpet of the fourth images in Figure 5 (b). This validates our idea that leverages the pre-trained T2I diffusion model as the prior for better generalization.
@@ 4.4 Intrinsic Image Decomposition
Intrinsic image decomposition [6] recovers albedo and shading properties from RGB images. It facilitates applications such as recoloring [5] and relighting [66]. Similar to PIE-Net [14], we use the mean square error $\mathrm{MSE}$ as the evaluation metric to compute the distance between the predicted and pseudo ground-truth estimations. We showcase the qualitative results in Figure 4 and quantitative comparisons in Table 3. The reported errors of most methods are an order of magnitude larger than ours. Notably, while our method and SPADE work well in both in- and out-of-domain settings, we found our method less influenced by artifacts in pseudo ground truth and to offer more reasonable estimations. More details are provided in the supplementary document.
@@
Figure 8: Single vs. multiple diffusion steps. In the single-step approach, we train the U-Net network to directly predict outputs from input images. The single-step approach does not generalize to unseen images well.
@@ 4.5 Ablation Study
Parameterization. We study the effect of employing various parameterizations. Specifically, we fine-tune the U-Net model to make different predictions in each reverse diffusion step for obtaining the final prediction: 1) predicting the input image $x$, predicting the output $y$ (similar to $x_{0}$-prediction in standard diffusion models), and v-prediction described in Eq. (3). We formulate each parameterization in detail in the supplementary document. The quantitative results are shown in Table 4 and the qualitative comparisons are presented in Figure 6. Predicting the input image $x$ generates accurate results in the in-domain setting. However, it fails to generalize to unseen domains. On the other hand, predicting the output $y$ demonstrates preferred generalization capability, but tends to produce blurry results with few details. We choose to use the v-prediction approach as it produces accurate results of arbitrary images.
Number of diffusion steps. We analyze the performance of our method with different numbers of diffusion steps in the inference stage. The quantitative comparisons on the normal prediction are presented in Figure 7. Using $5$ steps strikes a good balance between inference speed and accuracy. Furthermore, we study an extreme case of using a single step. We train the U-net to directly predict the output $y$ from the input image $x$. Although it shows competitive performance in the in-domain setting according to Figure 7, we find that the performance degrades significantly with the arbitrary input images, as examples shown in Figure 8. Considering the generalizability, inference speed, and accuracy, we use $5$ generation steps for all the other experiments.
@@ 5 Conclusion
In this work, we leverage a pre-trained T2I diffusion model for generalizable dense prediction. The core of our method is the design of the deterministic diffusion process that adapts the stochastic T2I framework for the deterministic prediction tasks. With low-rank approximation, the proposed approach learns the target tasks while retaining the inherent generalization ability of the T2I model. We show that with only a small number of labeled training data in a limited domain (i.e., $10$K bedroom images), our DMP scheme makes faithful predictions on arbitrary images. We believe that this work establishes a foundation for achieving ultimately-generalizable visual understanding in the future.
Appendix A Parametrizations
As described in Section 3.2, we empirically find that parametrizing the U-Net model through estimating v-prediction [62] performs favorably against predicting inputs or outputs. We detail the formulation of predicting inputs and outputs as follows. The U-Net model $\hat{x}_{\theta}$ predicting inputs is fine-tuned with the mean square loss:
$L=\mathbb{E}_{(x,y),t}\big{[}\|x-\hat{x}_{\theta}(y_{t},t)\|^{2}_{2}\big{]},$ (5)
and the reverse diffusion process is formulated as
$y_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\left(\frac{y_{t}-\sqrt{1-\bar{\alpha}_{t}}\hat{x}_{\theta}(y_{t},t)}{\sqrt{\bar{\alpha}_{t}}}\right)$ (6)
[…]
The U-Net model $\hat{y}_{\theta}$ predicting outputs is optimized with the loss function:
$L=\mathbb{E}_{(x,y),t}\big{[}\|y-\hat{y}_{\theta}(y_{t},t)\|^{2}_{2}\big{]},$ (7)
and the reverse diffusion process is
$y_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\hat{y}_{\theta}(y_{t},t)+\sqrt{1-\bar{\alpha}_{t-1}}x$ (8)
$t$ $=[T,\cdots,1],$
Appendix B Additional Experimental Results
B.1 Reliability of Off-the-Shelf Estimators
We indicate that the off-the-shelf estimators are not always reliable, especially the approach for intrinsic image decomposition. We demonstrate with the example of albedo estimation in Figure 9 that the off-the-shelf approach generates apparent artifacts in shadow areas, such as corners or floors under beds. The approach fails to recover the correct albedo but instead generates black patches. Consequently, SPADE also learns this pattern, but our model tends to correct artifacts by performing accurate estimation, manifesting the ability of generalization.
Figure 9: Qualitative comparisons on albedo estimation. SPADE [48] and the off-the-shelf approaches generate artifacts in dark areas.
B.2 Real-World Evaluation
NYU Depth v2.
Following Ke et al. [34], we evaluate our method on NYU Depth v2 [67] according to the protocol of affine-invariant depth evaluation [54]. We generate prompts with BLIP-2 [37] to use the model trained on synthetic bedroom images. We scale and shift the depth predictions to align ground truths by solving least-square fitting. The comparison against other approaches is shown in Table 5. DMP performs comparably with some previous methods trained with large-scale data.
Table 5: Comparison of performance on NYU Depth v2 [67].
# Training Samples NYU v2
[…]
MiDaS [54] 2M 11.1 88.5
Omnidata [16] 11.9M 310K 7.4 94.5
DPT [53] 1.2M 188K 9.8 90.3
Painter [78] 24K 8.0 95.0
Marigold [34] 74K 5.5 96.4
DMP 10K 12.0 86.5
ADE20K.
We also investigate the performance of semantic segmentation with more classes, e.g. 150 classes in ADE20K [89]. We follow the encoding strategy proposed by Wang et al. [78] and convert class indices into 3-digit numbers with a $b$-base system, which can be represented in the RGB space. However, the performance is unsatisfying (lower than $20\%$ accuracy). With the number of classes increasing, the differences between colors are less distinguishable. The unlabeled areas, which can be simply ignored when calculating loss in the image space, become hard to tackle in the latent space. We leave the application of real-world many-class semantic segmentation for future exploration.
We conduct another analysis with a subset of ADE20K containing only images with beds. The train-test split is constructed by applying the same filtering to the original splits, resulting in 1825 training images and 189 test images. We also generate prompts with BLIP-2. The results are presented in Table 6. The performance across real and synthetic domains is similar, especially for large items.
Table 6: Comparison of performance between the subset of ADE20K [89] and the synthesized bedrooms.
Bed Pillow Lamp Window Painting Mean
[…]
B.3 Additional Ablation Study
Modeling.
We vary the trainable layers and the presence of text conditions when fine-tuning the model. Since the example tasks we choose in this work are not directly conditional on text, providing text descriptions might not be necessary. Accordingly, training cross-attention layers is optional. Table 7 shows that text condition and training cross-attention layers help improve the performance of in-domain samples, but the difference in out-of-domain samples is unnoticeable between the settings. This result suggests that we can adopt curated real ground truth datasets without text descriptions for training at the expense of a subtle performance drop. Alternatively, we can generate prompts with image captioning [37], which may lead to better performance.
Table 7: Analysis of training cross-attention layers and providing text condition. Both improve the performance of in-domain samples but make little difference in out-of-domain data.
In-domain Out-of-domain
[…]
Self-attn 0.0606 0.1290 0.0890 0.1871
Self-attn + text 0.0605 0.1293 0.0876 0.1844
All attn + text 0.0514 0.1156 0.0872 0.1886
Size of Training Data.
We analyze the effect of varying the size of training data. We compare fine-tuning the model for normal estimation with $100$, $1$K, $10$K, and $100$K generated bedroom images. As shown in Figure 10, increasing data size over $10$K improves little performance, so we conduct the other experiments with $10$K training images.
Figure 10: Quantitative performance of normal estimation with different sizes of training data.
Table 8: NYU Depth v2 [67] performance comparison of models trained with real and pseudo ground truths.
[…]
Hypersim [56] Real 13.0 85.0
Bedrooms Pseudo 12.0 86.5
Quality of Training Data.
We examine the influence of data quality by comparing the models trained with real and pseudo ground truth. We use Hypersim [56] as the real ground truth and evaluate the models with NYU Depth v2 [67]. Table 8 shows that there is no significant difference between the two models. The model trained with pseudo ground truth even performs slightly better. We speculate that the data diversity may be an important factor. While Hypersim contains more than 70K images, the images are collected from only 461 scenes. Many of them are variations of camera views and distances. In contrast, the synthetic images, while all of them are bedrooms, are all distinct scenes, which present diverse compositions of objects.
Blending Inputs and Outputs.
IADB [23] proposes a deterministic framework where the diffusion process is formulated as a series of interpolations between observations and noise. Although their training strategy produces deterministic mapping of observations and noise, the correlation between observation and noise in each pair is stochastic due to unpaired sampling during training. We analyze the applicability of this framework to deterministic dense prediction problems by sampling paired inputs and outputs and fine-tuning from a pre-trained T2I diffusion model. With such adaptation, the differences between their framework and our approach are only the variance schedule and parametrization, where the importance weight of inputs linearly rises through the diffusion process, and the U-Net predicts $y-x$.
Table 9 shows the comparison between DMP and IADB on surface normal estimation, and Table 10 is the result of depth estimation. Figure 11 demonstrates that the images generated by the model fine-tuned through IADB have noise and inaccurate predictions.
In addition to $\alpha$-blending used by DMP and IADB, we investigate the effect of an advanced blending strategy, Poission blending [51], which blends source and target images by solving a least-square fitting while reserving the gradient of source images. We assume image gradients are meaningful in the latent space. The diffusion process is viewed as increasing the intensity of the mask for selection editing. We adopt an off-the-shelf PyTorch implementation [4]. The performance on surface normal estimation is shown in Table 9, and the example outputs in Figure 11 show that the image quality is unsatisfying.
Table 9: Comparions with IADB [23] and Poission blending [51] on surface normal estimation.
In-domain Out-of-domain
[…]
IADB [23] 0.0675 0.1416 0.0974 0.2017
Poission [51] 0.0868 0.1888 0.1201 0.2623
DMP 0.0514 0.1156 0.0872 0.1886
Table 10: Comparions with IADB [23] on depth estimation.
In-domain Out-of-domain
[…]
IADB [23] 0.3099 0.4982 0.1165 0.5049 0.3132 0.1467
DMP 0.1072 0.8861 0.1020 0.2117 0.6395 0.1360
Figure 11: Qualitative comparisons between different blending frameworks.
B.4 Additional Comparison
Table 11: Quantitative comparisons with ControlNet [86] and Palette [60] on 3D property estimation and intrinsic image decomposition.
Normal Depth Albedo Shading
In Out In Out In Out In Out
[…]
ControlNet [86] 0.1021 0.2216 0.1862 0.4032 0.1739 0.7681 0.1287 0.4398 0.4004 0.2253 0.0302 0.0402 0.0265 0.0336
Palette [60] 0.1643 0.3642 0.1881 0.4160 0.6889 0.2626 0.3604 1.0535 0.2270 0.4203 0.0203 0.0199 0.0304 0.0260
DMP 0.0514 0.1156 0.0872 0.1886 0.1072 0.8861 0.0041 0.2117 0.6395 0.1360 0.0051 0.0064 0.1020 0.0070
Table 12: Quantitative comparisons with ControlNet [86] and Palette [60] on semantic segmentation.
Bed Pillow Lamp Window Painting Mean
[…]
ControlNet [86] 0.5215 0.4820 0.3540 0.1436 0.4275 0.2936 0.4999 0.4190 0.3823 0.3257 0.4370 0.3328
Palette [60] 0.0347 0.0329 0.0019 0.0018 0.0013 0.0012 0.0119 0.0119 0.0005 0.0005 0.0101 0.0097
ControlNet.
ControlNet [86] proposes a conditional text-to-image framework with additional control, such as edges or human poses, which constrains structures and layouts of output images. Since it is also an image-to-image generative model, we train it to take input images as control and output estimations. The performance of estimating 3D properties and intrinsic images is presented in Table 11, and the segmentation results are shown in Table 12. It demonstrates weaker generalizability than our approach.
In addition, we analyze the influence of varying initial noise in Figure 12. While the rough structures of the images are controlled by the input images, the initial noise alters the details of estimations. This variation is not tolerated for dense prediction.
Figure 12: Results of ControlNet with different initial noise. The outputs are not deterministic.
Palette.
Besides training GAN-based generative models from scratch and fine-tuning pre-trained diffusion models with the approaches listed in Section 4.1, we additionally include training an image-to-image diffusion model from scratch for comparison. Following the design of Palette [60], we expand the input layers of the U-Net to encode the concatenation of input and output images, with the U-Net parameterized to predict noise. The same autoencoder in the pre-trained diffusion model is also adopted. The performance is shown in Table 11 and Table 12, which indicates the inability of this approach to handle categorical label maps.
B.5 Improving Compared Methods
Since the results of GAN-based generative models consistently outperform diffusion-based models in our experiments, we seek performance enhancement for diffusion-based approaches. All experiments are conducted on in-domain surface normal estimation.
Figure 13: Qualitative comparisons of SDEdit with starting from different time steps. A trade-off exists between the effect of style transfer and content preservation.
Table 13: Quantitative comparisons on in-domain surface normal estimation between different timesteps where the generation process of SDEdit starts. The performance improves at the expense of deviation from input image contents.
[…]
SDEdit.
[…]
Figure 14: Qualitative comparisons between DDIB and DDIB with Plug-and-Play (PnP). The image contents are reserved but not consistent with accurate normals.
Table 14: Quantitative comparisons on in-domain surface normal estimation between DDIB and DDIB with Plug-and-Play (PnP). The feature injection regulates the generated contents while improving performance.
[…]
DDIB 0.1849 0.4210
DDIB + PnP 0.1652 0.3634
DDIB.
As presented in Section B.7, DDIB is capable of generating images that are likely sampled from output distributions, but the contents and geometry of output results are not consistent with input images. We explore the approach to content consistency by adopting feature constraints proposed by Plug-and-Play (PnP) [76] for image-to-image translation, which injects the self-attention and convolution features of input images into output images. As shown in Figure 14 and Table 14, the structures and contents of output images of DDIB with PnP constraints highly resemble the input images, but the estimated normals remain inaccurate despite better quantitative performance.
IP2P.
We analyze the expressiveness of inverted tokens by varying the number of training tokens in IP2P (learned). While the performance is slightly improved in one metric of quantitative evaluation, shown in Table 15, Figure 15 reveals that the differences between the estimated results are not significant.
Table 15: Quantitative comparisons on in-domain surface normal estimation between different training tokens of IP2P (learned). The increased number of tokens does not guarantee improved performance.
[…]
1 0.3550 0.7181
2 0.3470 0.7790
4 0.3274 0.6384
Figure 15: Qualitative comparisons of IP2P with different training tokens.
B.6 Failure Cases
We demonstrate some examples of failure cases in Figure 16 for surface normal estimation, Figure 17 for depth estimation, and Figure 18 for semantic segmentation, where off-the-shelf approaches might provide more accurate prediction than our method.
Table 16: Style templates, where {} is replaced by original prompts.
  • anime artwork, {} . anime style, key visual, vibrant, studio anime, highly detailed
  • concept art, {} . digital artwork, illustrative, painterly, matte painting, highly detailed
  • comic, {} . graphic illustration, comic art, graphic novel art, vibrant, highly detailed
  • neonpunk style, {} . cyberpunk, vaporwave, neon, vibes, vibrant, stunningly beautiful, crisp, detailed, sleek, ultramodern, magenta highlights, dark purple shadows, high contrast, cinematic, ultra-detailed, intricate, professional
    surrealist art, {} . dreamlike, mysterious, provocative, symbolic, intricate, detailed
  • abstract style, {} . non-representational, colors and shapes, expression of feelings, imaginative, highly detailed
  • art deco style, {} . geometric shapes, bold colors, luxurious, elegant, decorative, symmetrical, ornate, detailed
  • vaporwave style, {} . retro aesthetic, cyberpunk, vibrant, neon colors, vintage 80s and 90s style, highly detailed
Figure 16: Failure cases of surface normal estimation.
Figure 17: Failure cases of depth estimation.
Figure 18: Failure cases of semantic segmentation.
B.7 Results of Compared Methods
We show the example images generated by the compared methods listed in Section 4.1. The results of surface normal estimation are in Figure 19, with depths in Figure 20, albedo in Figure 21, shading in Figure 22, and semantic segmentation in Figure 23.
Figure 19: Qualitative results of compared methods on surface normal estimation.
Figure 20: Qualitative results of compared methods on depth estimation.
Figure 21: Qualitative results of compared methods on albedo estimation.
Figure 22: Qualitative results of compared methods on shading estimation.
Figure 23: Qualitative results of compared methods on semantic segmentation.
Appendix C Implementation Details
Model Architecture and Optimization.
We use Stable Diffusion 1.4 as the pre-trained text-to-image model and adapt it with rank $=4$ for LoRA. We fine-tune the model for $50$K steps with batch size $8$ and learning rate $0.0001$ with a cosine decay schedule. The training takes around 14 hours with a single NVIDIA RTX 3090.
Generating Images.
We generate the training and test images by first generating a set of prompts with a large language model. The prompt for the language model is a template adapted from pix2pix-zero [50], where different scene keywords are filled in. The template is
“Provide a caption for a photo of a scene. The caption should contain many adjectives, should describe colors, styles, lighting and materials in the photo, should be in English and should be no longer than 150 characters. Caption:”.
The placeholder scene is replaced by “bedroom” for training images and in-domain test images. To generate out-of-domain test images for estimating 3D properties and intrinsic images, it is replaced by uniform sampling from the keywords in LABEL:tab:scenes.
Out-of-domain test images for segmentation are synthesized by varying the image styles of in-domain test images, for semantic categories should remain the same across training and test images. The prompts regulating the styles are listed in Table 16 borrowed from an online post [1].
Appendix D Applications
Surface normals and depths facilitate many vision tasks. We show by the examples of 3D photo inpainting [65] that precise depths improve 3D reconstruction from 2D images. Compared to the default depth estimator [54], the resulting videos produced with the depth maps generated by our approach have more accurate depth relationships between the objects. Please refer to the project website for visual demonstrations.
Table 17: Scenes categories of out-of-domain images.
airlock airplane cabin airport terminal airport ticket counter
alcove amusement arcade anechoic chamber indoor apse
aquarium arcade archive armory
indoor arrival gate art gallery art school art studio
artists loft assembly line indoor athletic field attic
auditorium auto factory indoor auto mechanics auto showroom
backstage indoor badminton court baggage claim ball pit
ballroom indoor bank bank vault banquet hall
indoor baptistry bar barbershop barrack
basement indoor basketball court bathhouse bathroom
indoor batting cage indoor bazaar beauty salon bedchamber
bedroom beer hall belfry bell foundry
berth berth deck betting shop bicycle racks
bindery biology laboratory indoor bistro indoor bleachers
indoor bomb shelter bookbindery bookstore indoor booth
indoor bow window bowling alley box seat boxing ring
breakroom indoor brewery indoor brickyard burial chamber
indoor bus depot bus interior indoor bus station butchers shop
indoor cabin cafeteria call center candy store
canteen backseat car interior frontseat car interior cardroom
cargo container interior indoor carport indoor casino catacomb
indoor cathedral catwalk chapel checkout counter
cheese factory chemistry lab indoor chicken coop indoor chicken farm
childs room interior choir loft indoor church indoor circus tent
classroom clean room indoor clock tower indoor cloister
closet clothing store cockpit coffee shop
computer room conference center conference hall conference room
confessional control room indoor control tower indoor convenience store
corridor courtroom interior covered bridge crawl space
cybercafe indoor dairy dance school darkroom
day care center delicatessen dentists office department store
departure lounge indoor diner dining car dining hall
dining room discotheque distillery indoor doorway
dorm room dress shop dressing room indoor driving range
drugstore editing room electrical room elevated catwalk
interior elevator elevator lobby elevator shaft engine room
entrance hall indoor escalator exhibition hall fabric store
indoor factory fastfood restaurant indoor ferryboat indoor firing range
fishmarket interior fitting room indoor flea market indoor florist shop
food court indoor foundry funeral chapel funeral home
furnace room galley game room indoor garage
indoor general store indoor geodesic dome gift shop great hall
indoor greenhouse indoor gun deck gun store indoor gymnasium
hallway indoor hangar hardware store hat shop
hatchery hatchway hayloft hearth
home office home theater hospital room indoor hot tub
hotel breakfast area hotel room indoor hunting lodge ice cream parlor
indoor ice skating rink indoor incinerator indoor inn indoor jacuzzi
indoor jail jail cell jewelry shop jury box
indoor kennel kindergarden classroom indoor kiosk kitchen
kitchenette lab classroom indoor labyrinth landing
laundromat lavatory lecture room legislative chamber
indoor library indoor lido deck limousine interior indoor liquor store
living room lobby locker room loft
indoor lookout station indoor lumberyard machine shop indoor market
martial arts gym maternity ward mess hall mezzanine
military hospital mill mine indoor mini golf course
indoor monastery morgue indoor mosque indoor movie theater
indoor museum music store music studio natural history museum
newsroom indoor newsstand nightclub indoor nuclear power plant
nursery nursing home indoor observatory office
office cubicles indoor oil refinery operating room optician
orchestra pit interior organ loft orlop deck ossuary
indoor outhouse oyster bar packaging plant palace hall
pantry paper mill indoor parking garage parlor
particle accelerator indoor party tent pawnshop penalty box
perfume shop pet shop pharmacy physics laboratory
piano store pig farm indoor pilothouse pizzeria
indoor planetarium playroom indoor podium portrait studio
indoor power plant print shop promenade deck indoor pub
pulpit pump room indoor quonset hut reading room
reception recreation room indoor recycling plant refectory
repair shop restaurant restaurant kitchen indoor restroom
revolving door riding arena indoor roller skating rink rolling mill
sacristy sauna sawmill science museum
scriptorium security check point server room sewer
sewing room shipping room indoor shipyard shoe shop
indoor shopping mall shower shower room shrine
indoor skywalk sporting goods store squash court stable
indoor stage staircase indoor steam plant indoor steel mill
storage room storeroom food submarine interior subway interior
supermarket sushi bar indoor swimming pool indoor synagogue
tearoom teashop television room television studio
indoor tennis court indoor tent textile mill indoor procenium theater
indoor round theater indoor seats theater thriftshop throne room
ticket booth indoor ticket window indoor tobacco shop toyshop
indoor track trading floor train interior rail indoor tunnel
road indoor tunnel turkish bath utility room utility tunnel
van interior indoor velodrome ventilation shaft vestry
veterinarians office videostore indoor volleyball court voting booth
waiting room walk in freezer indoor warehouse indoor washhouse
indoor water treatment plant wet bar whispering gallery wig shop
window seat winery witness stand workroom
workshop indoor wrestling ring youth hostel basketball arena
football arena hockey arena performance arena rodeo arena
soccer arena home atrium public atrium bakery kitchen
bakery shop airplane cargo deck choir loft cloakroom booth
cloakroom library cubicle office cubicle home dinette
vehicle dinette elevator door freight elevator ferryboat cargo deck
fitting room organ loft establishment poolroom home poolroom
spa massage room spa mineral bath corridor in a subway station platform in a subway station
turnstiles in a subway station platform in a train station station in a train station barrel storage in a wine cellar
bottle storage in a wine cellar